OpenStudy (anonymous):

Select steps that could be used to solve the equation 1 + 3x = -x + 4. add x, subtract 1, divide by 4 add x, subtract 4, divide by 4 subtract 3x, subtract 4, divide by 4 subtract 3x, subtract 4, divide by -4 subtract 1, add x, divide by 4

OpenStudy (anonymous):

lets just solve it \[1 + 3x = -x + 4.\] add \(x\) to both sides to get \[1+4x=4\] subtract 1 to get \[4x=3\] divide by 4 to get \[x=\frac{3}{4}\]

OpenStudy (anonymous):

actually there are a few right answers there, not just the one i used

OpenStudy (anonymous):

\[1 + 3x = -x + 4\] subtract \(3x\) \[1=-4x+4\] subtract 4 \[-3=-4x\] divide by \(-4\) \[\frac{3}{4}=x\]

OpenStudy (anonymous):

you could also subtract 1, add x, divide by 4